An optimizing consumer currently spends money on many goods. The last pair of sunglasses she bought cost $30 and gave her 10 utils of satisfaction. If this same consumer spends $10 on a hat and is still optimizing utility, how many utils will she receive from the purchase?
Answer #1

The MU/P for the goods should be the same in order to maximize utility.

MU/P (sunglasses) = 10/30

Price of hat = $ 10

Utils received from a hat = 10/30 x 10 = 10/3 = 3.33 utils